How to find x in this?
 
x + 3 = 10 – 2x + 18
 
One who will reply fast will be appreciated

Grade:11

7 Answers

SAHIL
7 years ago
x + 3 = 10 – 2x + 18
step 1: Take terms which include x in one side and remaining at other
x + 2x = 10+ 18 – 3
3x= 28-3
3x = 25
x= 25/3
and this is the answer
THANKS
 
nivesh kesharwani
6 years ago
step:1 transpose the 2x to rhs(right hand side)
x+2x = 10+18-3
3x = 28-3
3x = 25= 25/3and  this is the correct step.

Vivek
6 years ago
x+3=10-2x+18 arrange the all x to one side and all the constant to otherside this will make a equation 3x=25So x= 25/3. Answer
Om
5 years ago
Dear student
x + 3 = 10 – 2x + 18
step 1: Take terms which include x in one side and remaining at other
x + 2x = 10+ 18 – 3
3x= 28-3
3x = 25
x=25/3                      [By bringing x to LHS]
I hope this answer helps you
Thanking you
A fellow student
Om
5 years ago
Dear student
x + 3 = 10 – 2x + 18
step 1: Take terms which include x in one side and remaining at other
x + 2x = 10+ 18 – 3
3x= 28-3           
3x = 25
x=25/3                      [By bringing x to LHS]
 
Thanking you
A fellow student
TADA POOJAN
5 years ago
Dear student do the following steps to.solve the problem 
x + 3 =10-2x+18
x+2x=10+18-3
3x=25
x=25/3
Regards ,
Poojan
Ritthy
5 years ago
x + 3 = 10 – 2x + 18
x+2x+3= 10+18
3x+3 = 28
3x = 28-3
3x = 25
X = 25/3
Hope this answers helps you...